OpenStudy (anonymous):

a(subscript(n+1)) = [na(sub(n))] / [n+2] First 5 terms. (I'm struggling to do it because the a(sub(n+1)) thing is super confusing me.)

OpenStudy (campbell_st):

so what you have is \[a_{n +1} = \frac{na_{n}}{(n + 2)}\]

OpenStudy (campbell_st):

so are you given an initial value of n or a

OpenStudy (anonymous):

\[a _{n+1} = \frac{ na _{n} }{ n+2 }\] oh yeah. sorry. a1 = 2.

OpenStudy (campbell_st):

because all ot means to me n = 1 \[a_{2} = \frac{a_{1}}{3}\] n = 2 \[a_{3} = \frac{2 \times a_{2}}{4} \] n = 3 \[a_{4} = \frac{3 \times a_{3}}{5}\] etc

OpenStudy (campbell_st):

so just subsitute a1into the 1st equation to find a2... and then you can proceed to find all the others.

OpenStudy (campbell_st):

so for n = 1 \[a_{2} = \frac{1 \times 2}{3}\] get a value n = 2 \[a_{3} = \frac{ 2 \times a_{2}}{4}\] and continue

OpenStudy (anonymous):

my answer is: a1=2. a2=2/3. a3=1/3. a4=1/5. a5=1/15.

OpenStudy (anonymous):

thank you.

OpenStudy (campbell_st):

it seems to make sense... well done...